OpenStudy (anonymous):

How do I find the Inverse function for x^2-8x+8 ? p.s This is for Algebra 2 I want help not just the answer please.

zepdrix (zepdrix):

So we have a function:\[\Large f(x)=x^2-8x+8\]We'll replace our function notation with y, since it's a little easier to work with in these types of problems.\[\Large y=x^2-8x+8\]To find the inverse, we have all of our x's and y's trade places,\[\Large x=y^2-8y+8\]And from here to find the inverse function, we solve for our new y. Hmm so it's going to be a tad tricky since we have different degrees of y. We'll need to complete the square on the y's. To complete the square, we take half of the b term and square it. In this problem we have \(\large b=-8\). Half of that is \(\large -4\), then squared is \(\large 16\). So our constant term needs to be \(\large 16\). So we'll need another \(\large 8\). Let's add \(\large 8\) to each side.\[\Large x+8=y^2-8y+16\]

zepdrix (zepdrix):

Completing the square on the right simplifies it down to,\[\Large x+8=(y-4)^2\]To further solve for y we would take the square root and do some math stuff.